OpenStudy (anonymous):

A highway engineer specifies that a certain section of the roadway covering a horizontal distance of 2 km should have an upgrade of 8%. Compute the change in elevation of this section of the road.

OpenStudy (anonymous):

 \[\frac{\theta}{90^\textrm{o}}=8\% \\ \:\\ \theta = 8\%\cdot90^\textrm{o}=7.2^\textrm{o} \\ \tan\theta = \frac{\textrm{opp}}{\textrm{adj}}=\frac{\textrm{opp}}{2~\textrm{km}} \\ \:\\ \begin{align} \textrm{opp}&=2~\textrm{km}\cdot\tan\theta = 2~\textrm{km}\cdot\tan7.2^\textrm{o} \\ &\approx0.252659~\textrm{km}=252.659~\textrm{m} \end{align}\]